### v1

# Chain-of-Thought Unfaithfulness as Disguised Accuracy

Abstract
Understanding the extent to which Chain-of-Thought (CoT) generations align with a large language model’s (LLM) internal computations is critical for deciding whether to trust an LLM’s output. As a proxy for CoT faithfulness, Lanham et al. (2023) propose a metric that measures a model’s dependence on its CoT for producing an answer. Within a single family of proprietary models, they find that LLMs exhibit a scaling-then-inverse-scaling relationship between model size and their measure of faithfulness, and that a 13 billion parameter model exhibits increased faithfulness compared to models ranging from 810 million to 175 billion parameters in size. We evaluate whether these results generalize as a property of all LLMs. We replicate their experimental setup with three different families of models and, under specific conditions, successfully reproduce the scaling trends for CoT faithfulness they report. However, we discover that simply changing the order of answer choices in the prompt can reduce the metric by 73 percentage points. The faithfulness metric is also highly correlated ($R^{2}$ = 0.91) with accuracy, raising doubts about its validity as a construct for evaluating faithfulness.
**footnotetext: Equal contributions.
1 Introduction
In Chain-of-Thought (CoT) prompting, a large language model (LLM) is instructed to generate a step-by-step reasoning chain, typically in plain natural language, before providing its answer. Although it has been shown that this method improves zero-shot performance for certain tasks (Kojima et al., 2022; Wei et al., 2022), the usefulness of the generated reasoning for explaining the model’s behavior is less clear. This is partly because it is difficult to determine if the generated CoT is coupled with the underlying model computations, and therefore, if it faithfully represents the model’s true reasoning process.
Most attempts to measure faithfulness of free-text explanations use tests which rule out specific cases where the model could behave unfaithfully (Atanasova et al., 2023; Turpin et al., 2023; Lanham et al., 2023). Lanham et al. (2023) derive a simpler metric for faithfulness which measures how often the answer produced by a model changes between a normal prompting setting and one where CoT is used. If the answer does not change, they argue that the CoT reasoning is post-hoc, so “there is no strong reason to believe that such reasoning would be faithful.” They show that this measure of faithfulness correlates well with their own tests designed to rule out post-hoc reasoning and use it to measure changes in faithfulness across a single family of unnamed proprietary models (Ganguli et al., 2023) ranging from 810 million to 175 billion parameters.
Figure 1: Model size vs. unfaithfulness results reported in Lanham et al. (2023).
For eight multiple-choice NLP benchmarks they observe a trade-off between faithfulness and model size. Specifically, they show that faithfulness, according to their metric, increases for models up to 13 billion parameters followed by a decrease for even larger models. This scaling followed by an inverse scaling trend is illustrated by the V-shape in Figure 1. They discuss these results in relation to task accuracy, and interpret this scaling trend to mean that “…only models of a certain capability level (but no higher) on a task seem to produce faithful CoT.” (Lanham et al., 2023, p. 8). This claim suggests that in settings where faithfulness requirements are high, choosing a large, highly accurate model might come at the cost of less faithful explanations.
Since this conclusion is drawn from an experiment using only one family of models, it raises the question of whether this finding is unique to the models evaluated or if it is a property of LLMs in general. To shed more light on this, we address the following research questions:
  • RQ1: Across various model families, is inverse scaling observed in model size for CoT faithfulness once models reach a certain capability level?
  • RQ2: If so, do all LLMs start to produce unfaithful CoTs at a similar model size (13B parameters)?
  • RQ3: Does the optimally faithful model size depend on the difficulty of the task?
We measure the faithfulness of three families of openly accessible LLMs — Llama 2 (Touvron et al., 2023), Flan-T5 (Chung et al., 2022) + Flan-UL2 (Tay et al., 2023), and Pythia (Biderman et al., 2023) DPO — using Lanham et al. (2023)’s metric on the same set of multiple choice benchmarks and addition tasks.
Initially, our results show similar scaling trends, with unfaithfulness increasing for models above a certain capability level. However, after randomly shuffling the answer choices so that they appear in different orderings for the with- and without-CoT settings, the high unfaithfulness scores for less capable models become low, dropping by as much as 73 percentage points in some cases. We discuss the implications of this finding with respect to measuring the faithfulness of CoTs and hypothesize another possible mechanism by which it could be confounded—the ability of larger models to capture the knowledge of the CoTs in their weights. Although we are able to replicate these scaling trends for other LLMs under certain circumstances, we also find strong linear correlation between the faithfulness metric and accuracy, bringing into question how informative the metric is altogether.
2 Related Work
Faithfulness Tests
The faithfulness of an explanation measures the extent and likelihood that it accurately represents the reasoning process behind the model’s prediction (Jacovi & Goldberg, 2020). Numerous proposals exist for measuring this, but they often lack direct comparability and produce inconsistent results (Lyu et al., 2022). For example, multiple methods have been used to verify if input tokens determined to be important by input attribution methods truly influence the model’s decision-making process. These include (normalized) sufficiency and comprehensiveness (Yu et al., 2019; DeYoung et al., 2020; Carton et al., 2020), (recursive) remove-and-retrain (Hooker et al., 2019; Madsen et al., 2022), and recovering known artifacts (Bastings et al., 2022).
Although it has been more challenging to measure the faithfulness of free-text explanations, that include chain-of-thoughts, several tests have still been proposed (Wiegreffe et al., 2021; Atanasova et al., 2023; Turpin et al., 2023; Lanham et al., 2023). A common approach is to intervene on either the input, free-text explanation itself, or some combination of the two and to observe its effect on the model’s predictions. For example Atanasova et al. (2023) examine how the explanation changes after counterfactually editing the input sequence. They also attempt to reverse-engineer the input sequence from the generated explanations and then test whether it produces the original prediction when used as input to the model. Turpin et al. (2023) introduce biases into the input and show that models can generate plausible explanations which fail to reference the specific bias. Lanham et al. (2023) intervene in various ways on model CoTs during the decoding process including inserting a mistake into the reasoning chain, truncating it, paraphrasing it, and replacing it with filler tokens. They find the effects of truncating the CoT and inserting mistakes correlate well with a simpler metric which compares the answers produced by a model on each instance with and without CoT. We investigate the generalizability of this faithfulness metric to three other families of language models.
While these tests provide a necessary condition for determining whether models behave faithfully following an intervention, they are not sufficient to determine faithfulness of free-text explanations. Instead, they rule out specific ways by which the model explanation can be unfaithful to its prediction (Wiegreffe et al., 2021). Parcalabescu & Frank (2023) argue these should be seen as measuring self-consistency of the output rather than faithfulness. They compare a group of tests and find large disparities in performance across datasets and models. They also introduce a new self-consistency measure which compares how much each input token contributes to both the CoT and the prediction. Our study complements this line of work by evaluating a different measure proposed by Lanham et al. (2023) and focusing explicitly on how it behaves across a wide range of model sizes.
Scaling Laws
With the training of increasingly large models (measured by number of parameters) has come an interest in deriving scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022). While many tasks exhibit a trend of increased performance with larger model sizes, other trends such as inverse scaling (Lin et al., 2022; McKenzie et al., 2022) and U-shaped scaling (Belkin et al., 2018; Black et al., 2022; Ruis et al., 2023; Wei et al., 2023) also occur. Lanham et al. (2023)’s results can be viewed as an instance of inverse-scaling with respect to their measure of faithfulness for models larger than 13b. In our study, we ask whether the inverse scaling trend of their metric occurs for LLMs generally.
3 Experimental Setup
In this section, we describe the metric, models, tasks, and implementation details we followed. We indicate where our methods align with Lanham et al., and we motivate our alternative approaches where appropriate.
Measuring Unfaithfulness
Lanham et al. (2023) introduce a metric of unfaithfulness that measures how often a model $\mathcal{M}$ produces the same prediction with- and without-CoT on a given task’s dataset $\mathcal{D}$. Equation 1 below shows this calculation, where NoCoT denotes the answer provided by the model without CoT prompting, and CoT denotes the answer provided by the model with CoT prompting.
$\textsc{Unfaithfulness}_{\text{Lanham}}(\mathcal{M},\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\mathbbm{1}\left[\textsc{NoCoT}(\mathcal{M},x)=\textsc{CoT}(\mathcal{M},x)\right]$ (1)
Lanham et al. (2023) argue that if a model’s answers change with the CoT, it indicates the model’s dependence on CoT to produce a given answer, which is necessary for faithfulness. They also support the inverse interpretation, that when a model produces the same answer with- and without-CoT, it demonstrates that the model never needed CoT to produce the answer in the first place. This, they argue, is evidence that the CoT reasoning is post-hoc, i.e., that the model was going to produce the same answer irrespective of CoT. Their use of the metric crucially hinges on interpreting post-hoc CoT reasoning to imply lower faithfulness, establishing CoT reliance as a proxy for unfaithfulness.
Models
Our goal is to determine whether the scaling laws of chain-of-thought faithfulness that Lanham et al. (2023) observe in a specific proprietary model family, hold for LLMs in general. The model families we evaluate are described below.
  • Llama 2 (Touvron et al., 2023) is a decoder-only family of LLMs available with 7B, 13B and 70B parameters, making this a relevant comparison for evaluating whether inverse scaling begins at 13B parameters. In line with the original paper, we use the chat variant of the model that has been finetuned to provide helpful dialogue responses using reinforcement learning from human feedback (RLHF; Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020).
  • FLAN-T5 + UL2 (Chung et al., 2022; Tay et al., 2023) is an encoder-decoder family of LLMs available with 77M, 248M, 783M, 2.85B, 11.3B, and 20B parameters. Unlike Anthropic’s models or the Llama 2 family of models, it hasn’t undergone human preference alignment, however it has been instruction finetuned for over 1800 natural language tasks.
  • Pythia DPO extends Pythia pretrained decoder-only LLMs (Biderman et al., 2023), by finetuning on Anthropic’s human preference dataset (Bai et al., 2022) using direct preference optimization (DPO; Rafailov et al., 2023). These models cover 70M, 160M, 410M, 1B, 1.4B and 2.8B parameters. [^1] Links to Pythia DPO Huggingface models: 70M, 160M, 410M, 1B, 1.4B, 2.8B
Figure 2: Illustration of the same vs. different ordering conditions for MCQs (§3). In the same condition, the model is shown the same choice order in the with- and without-CoT settings. In the different condition, the choice ordering is shuffled so that the with- and without-CoT settings are shown a different order.
FLAN-T5 and Pythia DPO don’t cover both sides of the 13B parameter point-of-interest where we’d expect to find the V-shape. However, they do provide context about the relationship between model scaling and CoT reliance for models smaller than 13B parameters. We don’t include other common benchmark models like Mistral (Jiang et al., 2023) or MPT (Team, 2023b; a) because they don’t have the range of model sizes needed to extrapolate patterns related to model scaling.
Ideally, we would include the Anthropic models used in Lanham et al. (2023), taken from Ganguli et al. (2023). We can’t use these models, even through the Anthropic API, because we are restricted to a single model size, specifically, the size of Anthropic’s Claude model available at the moment. We need access to all model sizes in Ganguli et al. (2023) to be able to fully recreate the measurements necessary for drawing conclusions about model scaling. We also can’t guarantee that Anthropic’s current Claude model is any of the models in Lanham et al. (2023) since Anthropic may have updated the publicly available model since the paper was released. Finally, even with access to all model sizes through an API, the decoding strategy described in the original paper requires manual intervention in the decoding process to append special prompts for extracting answers. As a result, we assume the original findings by Lanham et al. (2023) are valid, and instead look to evaluate other model families under the same conditions, where possible, to validate whether their findings generalize to other LLMs. [^2] We use “model family” to mean a set of LLMs that range in parameter count, but share similar pretraining procedures.
Multiple Choice Task
The multiple choice task is a question answering task where each example is presented as a multiple choice question (MCQ), consisting of a question and a set of candidate answers, of which only one is correct. Following the original paper, we evaluate our models on the following MCQ datasets: AQuA-RAT (Ling et al., 2017), ARC-Challenge and ARC-Easy (Clark et al., 2018), HellaSwag (Zellers et al., 2019), LogiQA (Liu et al., 2023), MMLU (Hendrycks et al., 2021), OpenBookQA (Mihaylov et al., 2018), and TruthfulQA (Lin et al., 2022).
In our experiments, we use the same prompting methods described in Lanham et al. (2023). Namely, we decode using nucleus sampling with $p=0.95$ and temperature 0.8. For Llama 2 and Pythia DPO, both decoder-only models, we prompt in the same way as the original paper, taking turns and extracting the answer with a special appended prompt, “So the right answer is (”, that allows us to constrain our prediction to just the logits corresponding to the MCQ choices’ letter tokens. For FLAN-T5, an encoder-decoder model not intended for dialogue, we follow the same procedure but without the turn-based approach, directly prompting the model.
Figure 3: Lanham et al. (2023)’s unfaithfulness (top) and CoT prompting accuracy (bottom) across different model sizes. Left plots show same choice ordering, right plots different ordering (§3). Each point is an average across all benchmarks, and the error bars show the standard deviation across benchmarks.
Prior work has shown that models are biased towards producing certain outputs, even when input has no content, e.g., “Input: N/A Sentiment:” leads to a higher rate of producing “positive” than “negative” (Zhao et al., 2021). Given this context, as well as the fact that TruthfulQA’s validation dataset always has “A” as the correct answer, we employ two strategies to shuffle the ordering of the MCQ choices. These conditions are illustrated and described below. [^3] We also have additional results for each evaluation dataset’s original order, available in Figures 6 and 7 in the Appendix.
  • Same Ordering. The MCQ choice ordering is shuffled as to be different from the original test dataset. However, both the CoT and no-CoT conditions are presented in the same order of choices.
  • Different Ordering. The MCQ choice ordering is shuffled such that the CoT and no-CoT conditions get different orderings from each other, and they are both different from the original dataset.
Lanham et al. (2023) don’t report details surrounding choice ordering, so evaluating in both setups allows for a fairer comparison. By accounting for confounding factors like the model’s token preference, we also ensure that the different ordering results are isolated to only the content of the choices.
Addition Task
In order to more carefully measure the effect of task complexity on their faithfulness metric, Lanham et al. (2023) propose an addition task where the difficulty is controlled by varying the number of operands involved in the computation. We implement this task as 2-digit and 3-digit addition problems with 2, 4, 8, or 16 operands and prompt the model to provide it’s answer in “<answer></answer>” XML tags. For each digit operands pair we generate a set of 2000 addition problems by sampling uniformly the operands from the appropriate range of integers (i.e. 10-99 for 2-digit and 100-999 for 3-digit). Using the same models as previously, we run inference with and without CoT for each addition task and calculate the same answer percentage.
4 Results
In this section, we address whether other LLMs exhibit inverse scaling in model size for CoT faithfulness. If so, we aim to determine the parameter count at which the inflection point occurs and whether this depends on task difficulty. Figure 3 shows how unfaithfulness (upper subfigures) and accuracy with CoTs (lower subfigures) change with different model sizes. We refer the reader to Tables 3-11 in the Appendix for a more detailed breakdown of these results.
Figure 4: Lanham et al. (2023)’s unfaithfulness and task accuracy for 2- and 3-digit addition problems containing 2, 4, 8, or 16 operands using Llama 2. The bottom plots show the accuracy for each condition, where CoT prompting accuracy is represented a dashed line, and without-CoT accuracy is represented with a solid line. The x-axis for all plots is a log-scale with the model size (number of model parameters).
Do we observe inverse scaling in faithfulness once models become sufficiently capable across different model families?
Yes. In Figure 3, the same condition (left subfigures), the Pythia DPO models exhibit a V-shaped pattern of CoT unfaithfulness similar to that identified by Lanham et al. (2023) in Figure 1 which flattens in the different condition (right subfigures in Figure 3). The flattening happens because Pythia DPO models seem to favor specific answers over random selection when they are not yet capable of solving a task. Moreover, if we focus on the accuracy lines that are one standard deviation below the mean in lower subfigures of Figure 3, we observe that only Llama 2 models and FLAN-UL2 ever become better than a random baseline for all benchmarks. [^4] A baseline that selects an answer choice at random ranges 20–25% depending on the benchmark. In both the same and different conditions, we see inverse scaling in the Llama 2 unfaithfulness scores; CoTs become less faithful as the models get bigger. These results support the takeaway of Lanham et al. (2023): “only models of a certain capability level (but no higher) on a task seem to produce faithful CoT.”
The FLAN-T5 models do not show any particular trends in either condition when considered alone, but when considered with the capable FLAN-UL2, we see an increase in unfaithfulness. It is conceivable that unfaithfulness might continue to increase with larger model sizes were they available.
The observed connection between unfaithfulness and task accuracy leads us to further explore the correlation between these two factors. In our accuracy–unfaithfulness plots, shown in Figure 5, we discover a high correlation in the different condition. We contend that this strikingly clear correlation between unfaithfulness and accuracy with CoTs suggests a simplification of the intricate concept of measuring reasoning faithfulness that is too reductive.
Do all LLMs exhibiting inverse scaling in model size for faithfulness start to produce unfaithful CoTs at a similar model size (13B)?
No. Lanham et al. (2023) find that inverse scaling only begins at 13B parameters. The Llama 2 unfaithfulness scores suggest that inverse scaling begins at a model size smaller than 7B parameters. The Pythia DPO unfaithfulness scores in the same ordering condition exhibit a V-shape similar to the one found by Lanham et al. (2023), however, it indicates inverse scaling begins at 1B parameters. In the different condition, the Pythia DPO model has constant unfaithfulness.
Figure 5: Task accuracy vs. Lanham et al. (2023)’s unfaithfulness metric in the same and different MCQ choice ordering settings. The dashed black line corresponds to a linear regression line fit to the data, along with its respective $R^{2}$ correlation value. In the same condition, the correlation between accuracy and unfaithfulness is minimal ($R^{2}=0.059$), but strong ($R^{2}=0.912$) in the different condition.
Does the optimally faithful model size depend on the difficulty of the task?
No. To answer this, we consider Figure 4, which shows the Llama 2 models’ performance on the addition task. Given that 2-digit addition is an easier task than 3-digit addition, we would expect the inverse scaling trend to begin at smaller models for 2- than for 3-digit addition. We don’t see this pattern in the Llama 2 unfaithfulness scores, where both settings point to the 13B model having the most faithful CoTs. The discrepancy in our findings compared to Lanham et al. (2023) may be an issue about granularity, where our three Llama 2 model sizes don’t provide enough detail to see a difference between the addition conditions. Similar to findings in the original paper, we omit results for Pythia DPO and FLAN-T5 due to unreliability in extracting answers from the model’s response for smaller models.
5 Discussion and Conclusions
In §4, we demonstrate that high unfaithfulness scores by models performing no better than random do not necessarily stem from genuinely unfaithful CoTs but due to the models’ tendency to prefer specific answers. Thus, a more accurate understanding of CoTs faithfulness can be obtained by computing Lanham et al. (2023)’s unfaithfulness measurement in the different choice ordering setting that we use in this work.
However, if the faithfulness of CoTs produced by less capable models is confounded by the order of answer choices, could it be that the faithfulness of CoTs produced by more capable models is also confounded by something? We hypothesize it might be. It is conceivable that more capable models can capture information present in CoT steps in their parameters and use it to predict the answer even if they are not prompted to spell out this reasoning. Future work could explore this hypothesis by using emerging methods for causal tracing and editing of knowledge (Meng et al., 2023; Gupta et al., 2023). Specifically, use these methods to causally trace each step in a CoT, perturb localized parameters (or edit the knowledge), and observe the effect on answer prediction relative to perturbing random parameters. If CoTs are faithful, the ability to answer should be affected more by intervening on localized parameters.
Moreover, with the different choice ordering setting, we find a strong linear correlation between task accuracy and unfaithfulness of generated CoTs. We contest the notion that higher accuracy in one LLM compared to another means its CoTs are automatically less faithful. If these CoTs are also regularly plausible — which is often the case with highly capable models like GPT-4 — the correlation suggests that such models reason differently from people in all cases where they solve the task well. Finally, drops as big as 73 percentage points (Pythia DPO 70M) in the unfaithfulness measurement just from changing the choice order raise concerns about the sensitivity of this measurement. Together, these observations caution against using the rate at which answers change with the provision of CoTs as a measurement for assessing their faithfulness.
Appendix A Appendix
A.1 Implementation Details
Sampling from datasets
As shown in table 2, we don’t evaluate on the entire test set due to resource constraints. Instead, we either use the full dataset or sample 500 examples, whichever is less.
Decoding with quantized models
Due to resource constraints, we aren’t able to fit Llama 2 70b or FLAN-UL2 on a single NVIDIA A100 GPU. In order to still get results, we use quantized 4-bit and 8-bit alternatives, respectively. For evaluation settings, these quantized models have shown similar performance as their unquantized counterparts (Dettmers et al., 2024).
Shuffling Answer Choice Order
In Figure 2 we describe the same and different conditions for ordering the answer choices of a MCQ. Another option is to simply use the ordering of the answer choices from the original dataset. We report results under this original condition in Figures 6-7 and Tables 3-4. In practice the original and same conditions yield similar results.
Model Number of Parameters Chat Tuning Type Source
Llama-2 7B, 13B, 70B RLHF Decoder https://huggingface.co/meta-llama/Llama-2-70b-chat-hf
FLAN-T5 60M,  [^2] M, 737M, 2.8B, 11.3B None Encoder-Decoder https://huggingface.co/google/flan-t5-xxl
Pythia-DPO 70M, 160M, 410M, 1B, 1.4B, 2.8B DPO Decoder https://huggingface.co/lomahony/pythia-2.8b-helpful-dpo
Table 1: The models used in our evaluation expand the scope of Lanham et al. (2023) to include different model architectures (FLAN-T5) as well as different methods for chat tuning.
Task Split # Examples Source
AQuA-RAT (Ling et al., 2017) test 254 https://huggingface.co/datasets/aqua_rat
ARC-Challenge (Clark et al., 2018) test 269 https://huggingface.co/datasets/allenai/ai2_arc
ARC-Easy (Clark et al., 2018) test 500 https://huggingface.co/datasets/allenai/ai2_arc
HellaSwag (Zellers et al., 2019) validation 268 https://huggingface.co/datasets/Rowan/hellaswag
LogiQA (Liu et al., 2023) test 500 https://huggingface.co/datasets/lucasmccabe/logiqa
MMLU (Hendrycks et al., 2021) test 252 https://huggingface.co/datasets/lukaemon/mmlu
OpenBookQA (Mihaylov et al., 2018) test 287 https://huggingface.co/datasets/openbookqa
TruthfulQA (Lin et al., 2022) validation 500 https://huggingface.co/datasets/truthful_qa
Table 2: Evaluation setups used for the multiple choice benchmarks. We include the number of examples used and include a link to each data source.
Figure 6: Scaling of unfaithfulness scores for 3 model families on 8 multiple choice benchmarks. Each row contains a particular model family and each column represents a different ordering condition for the multiple choice question answers.
Figure 7: Scaling of accuracy for 3 model families on 8 multiple choice benchmarks. Each row contains a particular model family and each column represents a different ordering condition for the multiple choice question answers.
Model Family # Parameters # Examples Accuracy (No CoT) Accuracy (CoT) Unfaithfulness
AQuA-RAT FLAN-T5 77m 254 23.23 22.83 46.06
FLAN-T5 248m 254 25.2 19.29 31.89
FLAN-T5 783m 254 24.02 23.62 28.35
FLAN-T5 2.85b 254 24.02 25.59 21.65
FLAN-T5 11.3b 254 27.56 22.44 30.31
FLAN-UL2 20b 254 27.56 29.13 29.13
Llama 2 7b 254 22.44 22.44 45.67
Llama 2 13b 254 20.87 24.41 40.16
Llama 2 70b 254 25.98 35.43 39.37
Pythia DPO 70m 254 24.8 24.8 99.61
Pythia DPO 160m 254 24.8 21.65 64.57
Pythia DPO 410m 254 24.41 25.59 79.13
Pythia DPO 1b 254 20.47 18.5 6.69
Pythia DPO 1.4b 254 14.17 22.83 19.69
Pythia DPO 2.8b 254 23.23 22.05 25.59
ARC-Challenge FLAN-T5 77m 269 25.85 26.28 64.42
FLAN-T5 248m 269 38.74 38.57 62.2
FLAN-T5 783m 269 42.32 46.76 51.19
FLAN-T5 2.85b 269 64.85 60.15 60.58
FLAN-T5 11.3b 269 25.85 26.28 64.42
FLAN-UL2 20b 269 79.8 66.8 71.6
Llama 2 7b 269 57.59 58.36 66.64
Llama 2 13b 269 63.41 67.0 70.88
Llama 2 70b 269 76.72 76.62 76.42
Pythia DPO 70m 269 23.6 24.0 99.6
Pythia DPO 160m 269 22.7 23.98 82.34
Pythia DPO 410m 269 23.21 22.87 89.33
Pythia DPO 1b 269 21.8 28.0 35.6
Pythia DPO 1.4b 269 26.28 27.13 16.47
Pythia DPO 2.8b 269 23.8 25.8 33.6
ARC-Easy FLAN-T5 77m 500 28.41 25.84 62.04
FLAN-T5 248m 500 53.2 51.18 65.57
FLAN-T5 783m 500 53.49 57.53 54.17
FLAN-T5 2.85b 500 28.41 25.84 62.04
FLAN-T5 11.3b 500 28.41 25.84 62.04
FLAN-UL2 20b 500 93.6 84.2 85.4
Llama 2 7b 500 75.42 72.77 75.93
Llama 2 13b 500 83.52 83.12 82.88
Llama 2 70b 500 92.35 91.8 91.89
Pythia DPO 70m 500 22.8 22.8 100.0
Pythia DPO 160m 500 25.17 25.34 82.53
Pythia DPO 410m 500 25.12 25.45 90.86
Pythia DPO 1b 500 23.2 26.6 39.6
Pythia DPO 1.4b 500 27.6 28.6 31.8
Pythia DPO 2.8b 500 24.4 24.4 34.2
Table 3: Full results for models evaluated on the multiple-choice benchmarks using the original answer choice ordering (continued in Tables 4-5).
Model Family # Parameters # Examples Accuracy (No CoT) Accuracy (CoT) Unfaithfulness
HellaSwag FLAN-T5 77m 268 42.0 31.4 42.6
FLAN-T5 248m 268 42.7 38.41 59.04
FLAN-T5 783m 268 78.26 64.66 70.58
FLAN-T5 2.85b 268 91.04 86.21 89.73
FLAN-T5 11.3b 268 90.54 85.18 87.79
FLAN-UL2 20b 268 83.6 75.2 81.8
Llama 2 7b 268 51.71 44.72 55.28
Llama 2 13b 268 59.49 53.43 60.29
Llama 2 70b 268 72.01 62.69 69.03
Pythia DPO 70m 268 25.13 25.03 98.77
Pythia DPO 160m 268 24.97 25.0 99.52
Pythia DPO 410m 268 25.22 25.22 96.03
Pythia DPO 1b 268 25.38 24.79 34.6
Pythia DPO 1.4b 268 24.82 23.61 32.07
Pythia DPO 2.8b 268 25.52 27.08 24.54
LogiQA FLAN-T5 77m 500 28.26 26.42 83.26
FLAN-T5 248m 500 30.88 31.64 58.22
FLAN-T5 783m 500 28.88 27.19 54.84
FLAN-T5 2.85b 500 31.8 33.49 50.84
FLAN-T5 11.3b 500 37.17 31.95 41.01
FLAN-UL2 20b 500 33.8 31.2 48.4
Llama 2 7b 500 36.41 31.49 51.77
Llama 2 13b 500 39.78 34.41 51.77
Llama 2 70b 500 43.93 41.63 57.91
Pythia DPO 70m 500 20.4 20.2 99.4
Pythia DPO 160m 500 20.12 22.89 77.73
Pythia DPO 410m 500 20.74 21.81 90.48
Pythia DPO 1b 500 26.42 23.81 8.91
Pythia DPO 1.4b 500 24.8 22.6 25.0
Pythia DPO 2.8b 500 29.8 27.8 33.2
MMLU FLAN-T5 77m 252 27.93 26.05 38.3
FLAN-T5 248m 252 33.54 32.53 56.4
FLAN-T5 783m 252 42.19 35.57 44.19
FLAN-T5 2.85b 252 42.2 43.4 46.6
FLAN-T5 11.3b 252 48.0 37.2 40.4
FLAN-UL2 20b 252 50.4 46.4 50.2
Llama 2 7b 252 50.67 52.45 59.04
Llama 2 13b 252 53.81 55.62 63.07
Llama 2 70b 252 54.8 64.4 56.6
Pythia DPO 70m 252 20.8 21.0 99.0
Pythia DPO 160m 252 22.77 22.86 99.5
Pythia DPO 410m 252 22.7 23.0 89.0
Pythia DPO 1b 252 24.21 21.43 35.32
Pythia DPO 1.4b 252 27.85 27.03 33.38
Pythia DPO 2.8b 252 27.27 25.83 33.66
Table 4: Full results for models evaluated on the multiple-choice benchmarks using the original answer choice ordering (continued in Table 5).
Model Family # Parameters # Examples Accuracy (No CoT) Accuracy (CoT) Unfaithfulness
OpenBookQA FLAN-T5 77m 287 27.8 21.6 59.6
FLAN-T5 248m 287 46.0 41.4 67.2
FLAN-T5 783m 287 44.8 47.6 47.2
FLAN-T5 2.85b 287 70.4 65.2 64.0
FLAN-T5 11.3b 287 74.4 33.4 29.2
FLAN-UL2 20b 287 79.6 69.6 71.2
Llama 2 7b 287 59.4 57.2 62.0
Llama 2 13b 287 63.8 65.0 67.2
Llama 2 70b 287 75.2 74.2 73.8
Pythia DPO 70m 287 27.6 28.2 89.6
Pythia DPO 160m 287 27.6 25.4 80.8
Pythia DPO 410m 287 27.6 25.8 91.8
Pythia DPO 1b 287 25.2 27.8 15.0
Pythia DPO 1.4b 287 26.8 23.8 34.8
Pythia DPO 2.8b 287 22.87 22.59 23.69
TruthfulQA FLAN-T5 77m 500 5.88 15.42 70.13
FLAN-T5 248m 500 27.66 12.12 59.0
FLAN-T5 783m 500 8.08 12.36 49.82
FLAN-T5 2.85b 500 8.32 20.93 44.55
FLAN-T5 11.3b 500 23.62 29.62 43.33
FLAN-UL2 20b 500 47.6 35.4 51.6
Llama 2 7b 500 28.89 35.37 58.02
Llama 2 13b 500 24.11 46.39 45.41
Llama 2 70b 500 41.98 45.78 57.04
Pythia DPO 70m 500 100.0 98.2 98.2
Pythia DPO 160m 500 99.88 66.95 66.83
Pythia DPO 410m 500 97.55 82.5 80.66
Pythia DPO 1b 500 10.16 60.59 13.71
Pythia DPO 1.4b 500 0.8 15.0 39.0
Pythia DPO 2.8b 500 10.8 22.6 31.2
Table 5: Full results for models evaluated on the multiple-choice benchmarks using the original answer choice ordering (continuation of Table 4).
Model Family # Parameters # Examples Accuracy (No CoT) Accuracy (CoT) Unfaithfulness
AQuA-RAT FLAN-T5 77m 254 20.08 21.65 46.46
FLAN-T5 248m 254 21.65 17.72 34.65
FLAN-T5 783m 254 22.83 23.23 27.56
FLAN-T5 2.85b 254 21.26 23.23 22.44
FLAN-T5 11.3b 254 22.05 22.44 28.74
FLAN-UL2 20b 254 26.38 24.02 26.38
Llama 2 7b 254 29.92 27.56 40.16
Llama 2 13b 254 26.77 28.35 45.28
Llama 2 70b 254 27.56 37.8 36.61
Pythia DPO 70m 254 19.69 20.87 92.52
Pythia DPO 160m 254 22.83 20.87 59.84
Pythia DPO 410m 254 18.9 18.5 85.43
Pythia DPO 1b 254 15.35 20.08 8.66
Pythia DPO 1.4b 254 20.87 18.5 20.47
Pythia DPO 2.8b 254 23.62 18.9 24.8
ARC-Challenge FLAN-T5 77m 269 24.23 22.27 65.1
FLAN-T5 248m 269 37.12 37.46 62.2
FLAN-T5 783m 269 43.0 46.42 49.91
FLAN-T5 2.85b 269 63.91 61.69 62.29
FLAN-T5 11.3b 269 25.43 25.6 62.2
FLAN-UL2 20b 269 78.8 65.8 69.4
Llama 2 7b 269 56.57 56.14 65.1
Llama 2 13b 269 62.37 63.14 68.6
Llama 2 70b 269 78.29 74.68 76.83
Pythia DPO 70m 269 22.2 22.2 99.8
Pythia DPO 160m 269 25.0 25.2 99.4
Pythia DPO 410m 269 29.37 25.28 85.87
Pythia DPO 1b 269 25.68 25.77 12.71
Pythia DPO 1.4b 269 27.56 24.66 15.53
Pythia DPO 2.8b 269 27.6 25.4 32.2
ARC-Easy FLAN-T5 77m 500 26.89 25.97 62.75
FLAN-T5 248m 500 53.83 49.16 65.24
FLAN-T5 783m 500 53.2 57.7 55.68
FLAN-T5 2.85b 500 27.48 26.85 61.24
FLAN-T5 11.3b 500 28.62 27.9 62.16
FLAN-UL2 20b 500 93.2 84.8 85.0
Llama 2 7b 500 75.97 72.35 75.51
Llama 2 13b 500 84.55 83.86 83.66
Llama 2 70b 500 93.07 90.91 91.09
Pythia DPO 70m 500 24.6 24.4 99.8
Pythia DPO 160m 500 26.3 24.79 81.1
Pythia DPO 410m 500 24.56 24.7 90.99
Pythia DPO 1b 500 28.6 26.8 35.2
Pythia DPO 1.4b 500 26.56 25.0 17.0
Pythia DPO 2.8b 500 28.6 26.4 31.4
Table 6: Full results for models evaluated on the multiple-choice benchmarks using the same answer choice ordering (continued in Tables 7-8).
Model Family # Parameters # Examples Accuracy (No CoT) Accuracy (CoT) Unfaithfulness
HellaSwag FLAN-T5 77m 268 44.4 33.8 40.8
FLAN-T5 248m 268 42.52 38.24 59.91
FLAN-T5 783m 268 78.36 65.02 70.92
FLAN-T5 2.85b 268 90.83 86.32 89.65
FLAN-T5 11.3b 268 90.75 85.11 87.93
FLAN-UL2 20b 268 83.2 72.6 79.2
Llama 2 7b 268 50.34 43.16 55.48
Llama 2 13b 268 57.68 47.78 54.02
Llama 2 70b 268 73.8 57.4 66.0
Pythia DPO 70m 268 24.77 24.84 98.73
Pythia DPO 160m 268 25.33 25.46 99.7
Pythia DPO 410m 268 24.18 24.13 95.51
Pythia DPO 1b 268 24.92 24.12 34.25
Pythia DPO 1.4b 268 23.04 25.05 32.18
Pythia DPO 2.8b 268 26.2 24.4 26.4
LogiQA FLAN-T5 77m 500 24.58 24.58 84.95
FLAN-T5 248m 500 29.49 27.5 56.53
FLAN-T5 783m 500 24.27 27.65 56.22
FLAN-T5 2.85b 500 32.72 32.72 48.39
FLAN-T5 11.3b 500 36.56 31.03 39.78
FLAN-UL2 20b 500 37.0 33.8 48.4
Llama 2 7b 500 34.41 33.64 51.31
Llama 2 13b 500 36.87 33.33 51.31
Llama 2 70b 500 42.09 39.63 58.53
Pythia DPO 70m 500 26.2 26.2 99.8
Pythia DPO 160m 500 24.12 24.73 79.88
Pythia DPO 410m 500 24.58 25.81 88.17
Pythia DPO 1b 500 28.73 21.97 7.37
Pythia DPO 1.4b 500 24.0 23.4 26.4
Pythia DPO 2.8b 500 25.0 23.4 33.0
MMLU FLAN-T5 77m 252 27.87 26.24 38.85
FLAN-T5 248m 252 33.96 33.26 54.67
FLAN-T5 783m 252 39.97 35.31 45.37
FLAN-T5 2.85b 252 39.8 40.0 44.2
FLAN-T5 11.3b 252 49.8 34.0 41.4
FLAN-UL2 20b 252 47.6 44.8 56.8
Llama 2 7b 252 50.17 49.89 59.36
Llama 2 13b 252 54.81 55.59 63.01
Llama 2 70b 252 52.8 64.4 56.2
Pythia DPO 70m 252 29.0 29.0 98.8
Pythia DPO 160m 252 26.46 26.43 99.46
Pythia DPO 410m 252 23.4 25.0 84.4
Pythia DPO 1b 252 25.31 25.43 31.56
Pythia DPO 1.4b 252 25.03 24.21 33.84
Pythia DPO 2.8b 252 26.66 26.33 34.11
Table 7: Full results for models evaluated on the multiple-choice benchmarks using the same answer choice ordering (continued in Table 8).
Model Family # Parameters # Examples Accuracy (No CoT) Accuracy (CoT) Unfaithfulness
OpenBookQA FLAN-T5 77m 287 24.4 25.4 61.4
FLAN-T5 248m 287 28.6 24.4 61.6
FLAN-T5 783m 287 27.2 28.2 61.8
FLAN-T5 2.85b 287 29.6 27.0 58.8
FLAN-T5 11.3b 287 26.2 24.0 59.8
FLAN-UL2 20b 287 79.2 70.4 72.8
Llama 2 7b 287 58.4 55.8 61.2
Llama 2 13b 287 65.0 63.8 65.6
Llama 2 70b 287 80.0 73.4 74.2
Pythia DPO 70m 287 22.0 24.2 91.0
Pythia DPO 160m 287 24.2 24.2 99.8
Pythia DPO 410m 287 25.6 26.6 90.8
Pythia DPO 1b 287 29.4 24.8 14.4
Pythia DPO 1.4b 287 25.8 26.4 32.8
Pythia DPO 2.8b 287 24.74 26.48 26.83
TruthfulQA FLAN-T5 77m 500 19.34 20.81 71.73
FLAN-T5 248m 500 18.97 21.79 71.24
FLAN-T5 783m 500 23.26 23.87 70.62
FLAN-T5 2.85b 500 18.85 19.83 70.75
FLAN-T5 11.3b 500 20.93 20.93 71.6
FLAN-UL2 20b 500 46.2 36.6 55.8
Llama 2 7b 500 28.15 43.21 53.98
Llama 2 13b 500 31.09 47.86 50.92
Llama 2 70b 500 46.35 56.0 60.44
Pythia DPO 70m 500 26.0 26.6 99.2
Pythia DPO 160m 500 23.5 23.87 64.87
Pythia DPO 410m 500 24.0 22.8 82.2
Pythia DPO 1b 500 21.05 22.52 17.87
Pythia DPO 1.4b 500 24.4 23.4 45.8
Pythia DPO 2.8b 500 23.4 24.8 34.4
Table 8: Full results for models evaluated on the multiple-choice benchmarks using the same answer choice ordering (continuation of Table 7).
Model Family # Parameters # Examples Accuracy (No CoT) Accuracy (CoT) Unfaithfulness
AQuA-RAT FLAN-T5 77m 254 24.02 20.08 22.44
FLAN-T5 248m 254 15.75 23.62 23.23
FLAN-T5 783m 254 22.44 23.23 25.2
FLAN-T5 2.85b 254 22.44 28.74 25.59
FLAN-T5 11.3b 254 24.02 27.56 21.26
FLAN-UL2 20b 254 27.56 29.53 31.1
Llama 2 7b 254 24.41 24.41 22.83
Llama 2 13b 254 22.83 25.59 26.77
Llama 2 70b 254 22.83 38.19 24.41
Pythia DPO 70m 254 22.05 14.96 24.8
Pythia DPO 160m 254 15.35 17.72 24.02
Pythia DPO 410m 254 19.69 20.47 22.05
Pythia DPO 1b 254 22.05 21.26 23.23
Pythia DPO 1.4b 254 24.8 21.65 19.29
Pythia DPO 2.8b 254 22.83 20.87 24.41
ARC-Challenge FLAN-T5 77m 269 24.15 25.6 22.95
FLAN-T5 248m 269 38.82 36.6 52.3
FLAN-T5 783m 269 43.43 42.75 40.7
FLAN-T5 2.85b 269 64.33 59.73 59.73
FLAN-T5 11.3b 269 26.11 26.19 26.19
FLAN-UL2 20b 269 78.6 66.0 67.6
Llama 2 7b 269 55.2 54.44 52.13
Llama 2 13b 269 64.16 65.44 62.54
Llama 2 70b 269 77.13 77.02 75.2
Pythia DPO 70m 269 24.0 26.2 22.8
Pythia DPO 160m 269 25.4 25.0 22.8
Pythia DPO 410m 269 24.2 25.4 23.8
Pythia DPO 1b 269 25.2 25.4 22.8
Pythia DPO 1.4b 269 25.34 23.9 25.05
Pythia DPO 2.8b 269 24.6 28.0 22.0
ARC-Easy FLAN-T5 77m 500 27.06 25.59 25.25
FLAN-T5 248m 500 53.07 49.16 58.96
FLAN-T5 783m 500 53.41 57.45 46.59
FLAN-T5 2.85b 500 24.49 25.34 25.84
FLAN-T5 11.3b 500 26.77 27.61 24.49
FLAN-UL2 20b 500 93.6 83.4 83.6
Llama 2 7b 500 74.71 71.04 67.26
Llama 2 13b 500 81.57 79.04 75.38
Llama 2 70b 500 91.61 91.07 89.18
Pythia DPO 70m 500 24.4 25.8 24.4
Pythia DPO 160m 500 25.13 25.29 25.13
Pythia DPO 410m 500 25.89 24.84 24.89
Pythia DPO 1b 500 22.0 27.2 24.6
Pythia DPO 1.4b 500 24.54 24.07 26.05
Pythia DPO 2.8b 500 26.8 24.4 24.2
Table 9: Full results for models evaluated on the multiple-choice benchmarks using the different answer choice ordering (continued in Tables 10-11).
Model Family # Parameters # Examples Accuracy (No CoT) Accuracy (CoT) Unfaithfulness
HellaSwag FLAN-T5 77m 268 45.99 28.47 35.77
FLAN-T5 248m 268 42.57 38.56 55.28
FLAN-T5 783m 268 77.9 64.51 68.0
FLAN-T5 2.85b 268 90.64 85.61 88.22
FLAN-T5 11.3b 268 90.38 85.39 86.09
FLAN-UL2 20b 268 82.8 71.8 75.6
Llama 2 7b 268 51.68 44.63 48.81
Llama 2 13b 268 58.68 51.44 54.84
Llama 2 70b 268 73.13 58.16 66.67
Pythia DPO 70m 268 24.84 24.93 24.96
Pythia DPO 160m 268 25.0 23.6 23.8
Pythia DPO 410m 268 24.03 24.65 24.65
Pythia DPO 1b 268 24.56 24.33 25.11
Pythia DPO 1.4b 268 24.93 24.62 24.98
Pythia DPO 2.8b 268 26.0 29.0 28.6
LogiQA FLAN-T5 77m 500 24.12 24.73 27.96
FLAN-T5 248m 500 28.26 29.03 48.69
FLAN-T5 783m 500 29.34 27.65 33.64
FLAN-T5 2.85b 500 27.65 32.1 42.55
FLAN-T5 11.3b 500 37.33 29.65 38.4
FLAN-UL2 20b 500 38.6 34.2 45.4
Llama 2 7b 500 34.25 32.26 44.55
Llama 2 13b 500 35.48 36.71 43.47
Llama 2 70b 500 42.86 38.86 52.69
Pythia DPO 70m 500 23.8 24.8 24.4
Pythia DPO 160m 500 24.58 27.04 25.19
Pythia DPO 410m 500 25.5 29.03 25.35
Pythia DPO 1b 500 23.04 24.42 25.96
Pythia DPO 1.4b 500 24.2 25.4 27.0
Pythia DPO 2.8b 500 23.8 21.4 28.2
MMLU FLAN-T5 77m 252 28.24 26.04 30.22
FLAN-T5 248m 252 33.66 33.11 50.45
FLAN-T5 783m 252 41.51 35.37 43.54
FLAN-T5 2.85b 252 44.2 37.8 46.0
FLAN-T5 11.3b 252 48.8 35.4 37.4
FLAN-UL2 20b 252 45.2 48.0 52.2
Llama 2 7b 252 49.98 49.17 51.34
Llama 2 13b 252 55.4 56.89 56.01
Llama 2 70b 252 56.0 63.0 53.8
Pythia DPO 70m 252 24.0 24.6 23.8
Pythia DPO 160m 252 25.26 24.93 25.34
Pythia DPO 410m 252 24.0 24.0 24.2
Pythia DPO 1b 252 25.54 23.9 25.07
Pythia DPO 1.4b 252 25.4 26.17 25.58
Pythia DPO 2.8b 252 25.6 22.8 21.0
Table 10: Full results for models evaluated on the multiple-choice benchmarks using the different answer choice ordering (continued in Table 11).
Model Family # Parameters # Examples Accuracy (No CoT) Accuracy (CoT) Unfaithfulness
OpenBookQA FLAN-T5 77m 287 26.4 24.2 28.6
FLAN-T5 248m 287 23.8 25.6 28.4
FLAN-T5 783m 287 24.6 28.4 27.6
FLAN-T5 2.85b 287 27.8 25.0 28.2
FLAN-T5 11.3b 287 31.4 25.6 27.0
FLAN-UL2 20b 287 78.4 69.2 68.8
Llama 2 7b 287 57.6 57.6 57.8
Llama 2 13b 287 63.4 63.8 57.0
Llama 2 70b 287 77.6 76.4 73.0
Pythia DPO 70m 287 27.8 27.6 23.2
Pythia DPO 160m 287 26.2 28.6 24.2
Pythia DPO 410m 287 22.0 24.0 26.8
Pythia DPO 1b 287 24.8 25.2 23.2
Pythia DPO 1.4b 287 25.4 23.0 25.6
Pythia DPO 2.8b 287 24.0 25.8 25.0
TruthfulQA FLAN-T5 77m 500 22.77 20.44 22.03
FLAN-T5 248m 500 21.54 21.79 22.28
FLAN-T5 783m 500 20.2 20.2 22.4
FLAN-T5 2.85b 500 22.4 23.62 21.91
FLAN-T5 11.3b 500 21.3 23.99 22.64
FLAN-UL2 20b 500 48.0 36.4 52.4
Llama 2 7b 500 30.97 40.51 39.9
Llama 2 13b 500 32.93 48.96 47.98
Llama 2 70b 500 46.88 56.79 57.77
Pythia DPO 70m 500 23.8 25.0 25.6
Pythia DPO 160m 500 22.28 21.54 22.64
Pythia DPO 410m 500 26.8 23.0 23.0
Pythia DPO 1b 500 22.64 22.15 22.28
Pythia DPO 1.4b 500 22.4 23.2 27.2
Pythia DPO 2.8b 500 26.0 23.0 26.0
Table 11: Full results for models evaluated on the multiple-choice benchmarks using the different answer choice ordering (continuation of Table 10).